Examples for L^{A}TEX 2ε

Giuseppe Ghibò ⟨ghibo@galileo.polito.it⟩

October 9, 1994

SpecialHost is also compatible with L^{A}TEX 2ε. For instance the following commands

\begin{figure}[ht]
\fboxsep=3pt
\centering
\fbox{\includegraphics{pic2e.eps}}
\caption{A simple PostScript picture.} \label{fig:one}
\end{figure}

produce the picture shown in figure 1.

Figure 1: A simple PostScript picture.

Of course, we can do anything with the picture, for example apply it scaling or rotation. For instance with

\begin{figure}[ht]
\fboxsep=3pt
\centering
\fbox{\rotatebox{45}{\includegraphics{pic2e.eps}}}
\caption{A simple, rotated, PostScript picture.} \label{fig:two}
\end{figure}

we obtain the picture showed in figure 2. We may also combine either scaling and rotation; for example

\begin{figure}[ht]
\fboxsep=3pt
\centering
\fbox{%
\rotatebox{45}{%
   \scalebox{0.5}[1]{%
      \rotatebox{90}{%
         \includegraphics{pic2e.eps}%
      }
      \hspace{0.2cm}
      \includegraphics{pic2e.eps}

Figure 2: A simple, rotated, PostScript picture.

      }
      \hspace{0.2cm}
      \includegraphics{pic2e.eps}
   }
   \hspace{0.5cm}
   \fbox{\includegraphics{pic2e.eps}}
}
\caption{A simple, rotated, PostScript picture.}
\label{fig:three}
\end{figure}

which picture produced is shown in figure 3.

Unfortunately, TEX text cannot be rotated or scaled by SpecialHost...

Figure 3: A simple, rotated, scaled, …, PostScript picture. 3